A watermelon farmer is operating in a perfectly competitive market. The market price of watermelon is $5 per pound and each farmer produces 1,000 pounds per week. The average variable cost per unit is $3 per pound and the average fixed cost per unit is $1.
a. What is a watermelon farmer’s profit in the short run? Explain.
b. What happens to the total number of farms in the long run? Why?
c. If technology reduces the cost of watermelon farming while the demand for watermelon increases, what happens to price in the long run? Explain.

thank you.
Answer #1

a). Profits= Total revenue - Total cost.

The total revenue is the price multiplied the quantity sold, and the total cost is the sum of both the total fixed costs and the total variable costs. The total fixed cost is average fixed cost multiplied by the quantity and the variable cost is average variable cost multiplied by quantity.

Total revenue= Priceimesquantity.

= 5 × 1000

  5000

Total cost= TFC+TVC.

TFC= AFCimesQuantity.

= 1 × 1000

  = 1000

TVC= AVCimesQuantity.

  3 × 1000

  =3000

Total cost= 40

Total economic profit  

  5000-4000

= 1000.

b). Since the farmers are earning positive economic profit this would attract new producers and in the long run the number of producers will increase.

c). The technology would reduce the cost of production and this would enable the farmers to supply more watermelon into the market, here the same time the demand for the watermelon also had increased and this would be a simultaneous shift in the demand and the supply curves. So in the long run the price will remain the same.

In the long run the price is determined by the market supply and the market demand.